OpenStudy (deadpooi):

Solve the equation: -3 2/5 + b = 8 1/5 A. b = 11 3/5 B. b = 4 4/5 C. b = 5 3/5 D. b = 11 1/5

OpenStudy (mysticfury):

re order the equation 8 - -3 2/5 = b check your work by -3 2/5 + b = 8

OpenStudy (mysticfury):

sorry it was 8 1/5 , not just 8 what did you get for an answer when you re ordered it?

OpenStudy (deadpooi):

wait let me get what I got. vs your answer

OpenStudy (mysticfury):

8 and 1/5 - -3 and 2/5 = ? kk

alones (alones):

believe it was my brave act. −3 2/5+b=8 1/5 Simplify both sides of the equation. b+ −17/5 = 4/15 add 17/5 to both sides b+ −17/5 + 17/5=41/5 + 17/5 b=58/5 b=11 3/5 and so the answer is b=11 3/5

OpenStudy (deadpooi):

Wait

OpenStudy (deadpooi):

Oh wow I got b = 11 1/5

OpenStudy (mysticfury):

yep thats what i had @AloneS

OpenStudy (deadpooi):

Thank you for helping me.

OpenStudy (mysticfury):

not sure who your talking to but you get a your welcome from me